OpenStudy (anonymous):

Mrs. Santos invested P15, 000 in a money market fund for 2 months. The yearly rate of interest was 12%. How much did the money earn? What was the amount at the end of 2 months?

OpenStudy (anonymous):

total amount after 2 months=15300

OpenStudy (anonymous):

could you show me how did you solved this!?

OpenStudy (anonymous):

= (2/12)*0.12*15000 =300 ==> total =15000+300 =15300

OpenStudy (anonymous):

yes as shown by raheen

OpenStudy (dumbcow):

Simple interest A = P(1+i*t) A = 15,000(1+(.12)(1/6))

OpenStudy (anonymous):

12 month's interest=12% 2 " " =2% 15000*(102/100)=15300

OpenStudy (anonymous):

u get it??

OpenStudy (anonymous):

yup... thanks.

OpenStudy (nilankshi):

good

OpenStudy (anonymous):

what's good?

OpenStudy (nilankshi):

every thing is good

OpenStudy (anonymous):

oh.nice

OpenStudy (nilankshi):

thx